OpenStudy (anonymous):

Solve for x in the proportion x/2x+5=-5/x x=5 x=15 x=-25 x=-5

zepdrix (zepdrix):

You didn't use brackets, so the fraction on the left is a little hard to read. Is this what it's suppose to look like? I just wanna make sure first :) \[\large \frac{x}{2x+5}=-\frac{5}{x}\]

OpenStudy (anonymous):

yea exept the 5/x is negitive @zepdrix

zepdrix (zepdrix):

There's a negative in front of the one I wrote out, don't you see it silly?? c: We'll start by dealing with these fractions. Getting rid of the denominators. First, Multiply both sides by x,\[\large x\frac{x}{2x+5}=-\frac{5}{x}x\] Notice how on the right side, we are both multiplying and dividing by an \(\large x\). Those will cancel out for us,\[\large x\frac{x}{2x+5}=-\frac{5}{\cancel x}\cancel x \qquad \rightarrow \qquad x\frac{x}{2x+5}=-5\]We'll multiply the x's on the left side,\[\large \frac{x^2}{2x+5}=-5\] Ok so we've successfully dealt with the denominator on the right side! We're going to do the same for the other side. Is this making a little bit of sense? :O I hope

OpenStudy (anonymous):

x = 15 is that right? @zepdrix

zepdrix (zepdrix):

Hmm no :O Not so much

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

x = 5? @zepdrix

zepdrix (zepdrix):

yayy good job \c:/